OpenStudy (anonymous):

Solve this system. Please put your answer as an ordered pair. -x - 7y = -14 -4x - 14y = -28 i give medal for answer

OpenStudy (anonymous):

well you can multipli -x - 7y by -4 and ofcourse multipli -14 by -4 too ;) so it would be : 4x + 28y = 56

OpenStudy (anonymous):

so u have this system now : 4x + 28y = 56 -4x - 14y = -28 now add both sides : 4x -4x + 28y - 14y = 56 - 28

OpenStudy (anonymous):

and as 4x - 4x = 0 then, 28y - 14y = 56 - 28 now solve for y ;)